2. Sections 4.3,4.5,4.6 Graphing:Consider the function f (x) = sin(2x) + cos(2x)on the interval [0, 1]. For this question, giPlease show all work. Thank you!
Answer #1

f (w) Sin (2x) + cos (2x) [., a f(x) = sin(2x) + COS(22)) d (sin (2x)) + Šx (cos(2x)) Cos (2x) (2) + - Sin (2x) (2) f(x) 2 CoPut entrene points in tw)=sin(2x) + cos2x) Sin (215)) + Cos(2(T) = 2 x - T 8 t() x= 5T f(3) = sin(2)) + COS( 2(32)). - localPut Intlection points มา (x) = Sin (2x) + Cos (2x) 37 +(अ) = Sin(AP)) + as (2C))0 a ग (8) = 30()+ ces ((2)) = 0 Inthection po